OpenStudy (anonymous):

sin2xcosx-cos2xsinx =(sqrt3/2)

OpenStudy (anonymous):

\[\sin 2x \cos x-\cos 2x \sin x=\frac{ \sqrt{3} }{2 }\] \[\sin \left( 2x-x \right)=\sin60,\sin \left( 180-60 \right)\] x=?

OpenStudy (anonymous):

120

OpenStudy (anonymous):

sin x=sin 60,sin 120 x=60,120